A metal rod with a length of 30.0 cm lies in the xy-plane and makes an angle of 33.6 ∘ with the positive x-axis and an angle of 56.4 ∘ with the positive y-axis. The rod is moving in the +x-direction with a speed of 6.80 m/s . The rod is in a uniform magnetic field B⃗ =(0.100T)i^−(0.300T)j^−(0.0800T )k^.

What is the magnitude of the emf induced in the rod?
Answer #1

Eng - (VXB). v=6.81 L=0.3(26333.6i + 51336.j) 8 = 0.17 -0.3 -0.06€ vxBle boi (onli -o-sj -0.05). ixjak ixk -- => 6.8 x0.1 (ix